OpenStudy (anonymous):

divide 7w^2+3w-4 by w+1

OpenStudy (anonymous):

7w^2+3w-4=7w^2+7w-4w-4 =7w(w+1)-4(w+1) =(7w-4)(w+1) now ur question becomes:(7w-4)(w+1)/(w+1)=(7w-4)

OpenStudy (anonymous):

thats not what i asked. i asked how to divide them. idk how to do that

OpenStudy (anonymous):

That is exactly what you asked. A better question might be: I don't understand, can you explain, please?

OpenStudy (anonymous):

no its not i asked to divide them not factor them

OpenStudy (shadowfiend):

The point they are both making is that the most expedient way of dividing the two is to factor 7w^2 + 3w - 4 and then divide.

OpenStudy (shadowfiend):

Here's the long division version, however:

OpenStudy (anonymous):

that is what i want

OpenStudy (shadowfiend):

\[\begin{align} &7w\\ &- - - - - - -\\ w + 1 |& 7w^2 + 3w - 4 \\ & 7w^2 + 7w \\ & - - - - - \\ & 0 - 4w - 4 \end{align}\] \[\begin{align} &7w - 4\\ &- - - - - - -\\ w + 1 |& 7w^2 + 3w - 4 \\ & 7w^2 + 7w \\ & - - - - - \\ & 0 - 4w - 4 \\ & - 4w - 4 & - - - - - & 0 \end{align}\]

OpenStudy (shadowfiend):

Eh, kind of failed that last line there, sorry.

OpenStudy (shadowfiend):

Kenya, does the above make sense?

OpenStudy (anonymous):

ya i just kind of copied it i dont really undestand the whole dividing polynomials thing and i have a end-of-course on this on thursday

OpenStudy (shadowfiend):

Ok, well, the most straightforward way to see it is, above. You're dividing 7w^2 + 3w - 4 by w + 1. Now, w + 1 has two parts, the w, and the 1. That means you'll have to deal with only two parts of 7w^2 + 3w - 4 at a time. We start at the left, with 7w^2 + 3w. We concentrate on getting the w to turn into a 7w^2, which is why we use 7w. That gives us 7w(w + 1) = 7w^2 + 7w. Then we subtract that from the full polynomial, 7w^2 + 3w - 4. That leaves us with -7w - 4.

OpenStudy (shadowfiend):

Sorry, that leaves us with -4w - 4. Now that's another two parter, so we focus on -4w. We can turn w into -4w by multiplying by -4. So we, have -4(w + 1) = -4w - 4. That's the same as -4w - 4, so they cancel out, and we've successfully divided the two polynomials.

OpenStudy (anonymous):

wow that was like gibberish?????? to me sry

OpenStudy (shadowfiend):

Ok, so the point of dividing 7w^2 + 3w - 4 by w + 1 is to find something where, if we multiply it by w + 1, we get 7w^2 + 3w - 4. Right?

OpenStudy (anonymous):

ya

OpenStudy (shadowfiend):

Ok, so we start by seeing how we can get w to turn into 7w^2. We can do that by multiplying w by 7w.

OpenStudy (anonymous):

kk????

OpenStudy (shadowfiend):

Once we've found that we want to multiply by 7w, we multiply all of (w + 1) by it, and we get 7w^2 + 7w. Then, to reach the next step, we subtract it from the original polynomial, which is 7w^2 + 3w - 4. 7w^2 - 7w^2 = 0 3w - 7w = -4w - 4 - 0 = -4 So we end up with -4w - 4.

OpenStudy (shadowfiend):

Then, we need to find a way to turn w into -4w. We can do that by multiplying by -4.

OpenStudy (shadowfiend):

We do the same thing as above, multiply w + 1 by -4, which gives us -4w - 4, then subtract that from the remainder from the last subtraction, which was also -4w - 4: -4w - (-4w) = 0 -4 - (-4) = 0 And so we have 7w (the first part) + (-4) (the second part) as our result. This is the same as 7w - 4.

OpenStudy (anonymous):

k